On the day a person is born, a deposit of $50,000 is made in a trust fund that pays 8.75% in interest, compounded continuously.
A.) Find the balance on the person 35th birthday
B.) How much longer will the person have to wait for the balance to double?

Answer by rfer(16322)

A=50000(2.7183)^.0875*35
A=50000(2.7183)^3.0625
A=50000*21.38138
A=1,069,069.02
-------------------
Rule of 72
72/8.75=8.2 yrs
not exact but very close